Hello and thank you for being available.. Can you help me? Caan you simplfy or breakdown how to get there for me?
What are the two consecutive odd numbers of 56?

Answer by bucky(2097)

If x represents an odd number, then the next consecutive odd number is 2 more than x or x + 2.
{Think of 3 being an odd number. The next consecutive odd number is 5 and that is 2 more than 3.)
.
Add the two consecutive odd numbers together ... that is add x and x + 2 to get the sum 2x + 2.
.
The problem tells you that this sum is 56. In equation form this becomes:
.
2x + 2 = 56
.
Get rid of the 2 on the left side by subtracting 2 from both sides to get:
.
2x = 54
.
Solve for x by dividing both sides by 2 and the result is:
.
x = 54/2 = 27
.
So the first odd integer is 27 and the next consecutive odd integer is 2 more than that
or 29.
.
Check ... do 27 and 29 add up to 56? Yes, they do. Therefore, the answer to this problem is
that 27 and 29 are the two consecutive odd numbers that total 56.
.
Hope this helps you to understand the problem and how to solve it.
.